# Stacked Intelligent Metasurface-Enhanced Wideband Multiuser MIMO OFDM-IM Communications

### v1

@@ VI Conclusions
This paper proposed a SIM-enhanced multiuser OFDM-IM transceiver together with a physics-guided DUN tailored for wideband communications. On the system side, OFDM-IM is integrated as a structural complement to SIM: activating only a subset of tones widens the effective equalization bandwidth attainable by a practical $L$-layer SIM while simultaneously lowering the RF-output PAPR. On the objective side, we established a rigorous bridge from the worst-link BER to a distance-weighted per-tone SINR, thereby converting reliability optimization exactly into a max-min SINR problem with a single second-order cone programming backbone that enables joint phase-power-index design. On the solver side, we unfolded PGD into the proposed UPGD-Net, which enforces unit-modulus feasibility and monotone descent layer-wise and learns only a small set of per-iteration step sizes to support frame-rate reconfiguration under varying activation matrices.
Comprehensive simulations have corroborated these design choices. The proposed framework converges rapidly and stably, typically requiring 22\, $3\times$ fewer iterations than fixed-step PGD for comparable accuracy, and exhibits a clear layer-depth sweet spot. Compared with conventional OFDM-IM with digital ZF, the SIM-enhanced OFDM-IM achieves a consistent 22\, 3\,3 dB advantage at the target $\text{BER}\!=\!10^{-3}$, while its sum rate curve lies $80\%$ above the digital ZF counterpart across the tested SNR range. An activation–layout sweep further shows that the operating point (N,V)=(4,2)(N,V)\!=\!(4,2) strikes a favorable balance among BER, spectral efficiency, and detector complexity under perfect CSI. Finally, time-domain waveform analysis over the baselines confirms a reduced 22\, 5\,5 dB PAPR, underscoring the RF efficiency of the proposed architecture.
Beyond validating feasibility, this paper yields two design insights. First, structure design can broaden the bandwidth of analog processors: frequency-domain sparsity from OFDM-IM expands the usable equalization bandwidth of a finite-depth SIM. Second, the model-based network gains enhanced understanding capability of data: embedding exact PGD dynamics in a learnable scaffold (UPGD-Net) delivers near-optimal performance at frame-level latency while preserving interpretability and constraint satisfaction. These results highlight a practical path to reliability-aware, hardware-efficient wideband multiuser MIMO OFDM systems based on programmable metasurfaces.

### v2

@@ VI Conclusions
This paper proposed a SIM-enhanced multiuser OFDM-IM transceiver together with a physics-guided DUN tailored for wideband communications. On the system side, OFDM-IM is integrated as a structural complement to SIM: activating only a subset of tones widens the effective equalization bandwidth attainable by a practical $L$-layer SIM while simultaneously lowering the RF-output PAPR. On the objective side, we establish an analytical bridge from the worst-link BER to a distance-weighted per-tone SINR, thereby converting reliability optimization into a robust max-min SINR surrogate that enables phase optimization under dynamic index-power constraints. On the solver side, we unfold PGD into the proposed UPGD-Net, which enforces unit-modulus feasibility and monotone descent layer-wise and learns only a small set of per-iteration step sizes to support frame-rate reconfiguration under varying activation matrices.
Comprehensive simulations have corroborated these design choices. The proposed framework converges rapidly and stably, typically requiring 22\, $3\times$ fewer iterations than fixed-step PGD for comparable accuracy and exhibiting a clear layer-depth sweet spot. Compared with conventional OFDM-IM with digital ZF, the SIM-enhanced OFDM-IM achieves a consistent 22\, 3\,3,dB advantage at the target $\text{BER}\!=\!10^{-3}$. Crucially, its sum rate curve surpasses that of a large-aperture hybrid benchmark with $100$ antennas in the high-SNR regime, while offering decisive energy efficiency benefits due to the quasi-passive SIM structure. Furthermore, the proposed BER-driven design demonstrates strong resilience to channel estimation errors, maintaining effective beamforming where conventional zero-forcing fails. An activation–layout sweep further shows that the operating point (N,V)=(4,2)(N,V)\!=\!(4,2) strikes a favorable balance among BER, spectral efficiency, and detector complexity under perfect CSI. Finally, time-domain waveform analysis confirms a reduced 22\, 5\,5 dB PAPR, underscoring the RF efficiency of the proposed architecture.
Beyond validating feasibility, this paper yields two design insights. First, structure design can broaden the bandwidth of analog processors: frequency-domain sparsity from OFDM-IM expands the usable equalization bandwidth of a finite-depth SIM. Second, the model-based network gains enhanced understanding capability of data: embedding exact PGD dynamics in a learnable scaffold delivers near-optimal performance at frame-level latency while preserving interpretability and constraint satisfaction. These results highlight a practical path to reliability-aware, hardware-efficient wideband multiuser MIMO OFDM systems based on programmable metasurfaces.